OpenStudy (anonymous):

How do you find S(sub11) for1+2+4+8+...?

OpenStudy (valpey):

I think you are asking: "what is the sum of the first 11 terms". It is the 12th term minus 1.

OpenStudy (anonymous):

It is simple like"1+1 = 2" so it is not except you add that the summoned may not be repeated For example: N = { 1, 2, 3 ,...} is closed under addition Z = { ... , -2, -1, 0 , 1, 2 , ...} is closed under add and sub.

OpenStudy (anonymous):

hmmmm ok

OpenStudy (anonymous):

Thank you :p

OpenStudy (anonymous):

It's kinda like this "3 + 6 (mod 7) = 9 (mod 7) [9 ÷ 7 = 1 (quotient) ... 2 (remainder)] = 2 (mod 7) "r (mod d) (r = remainder, d = divisor)"